OpenStudy (anonymous):

A part of Php8,900 is invested at 6% and the remainder at 4%. The annual income from 4% investment exceeds the income from the 6% investment by Php74. Represent the amount of each investment

OpenStudy (shinalcantara):

let x be the amount invested @ 6% y be the amount invested @ 4% x+y = 8,900 ---eqn 1 0.04y = 0.06x +74 y = 1.5x + 1850 ----eqn 2 substitute eqn 2 to eqn 1 x + 1.5x +1850 = 8,900 2.5x = 8900 - 1850 x = 2,820 substitute to eqn 1 2,820 + y = 8,900 y = 8,900 - 2,820 y = 6,080

OpenStudy (anonymous):

THANK YOU SO MUCH @shinalcantara